# Modeling the precipitation distribution by radius and pore size during drying of an impregnated sphere

### v1

@@ 3.4 Distribution of precipitation by pore size
The capillary bundle model, which is used in the Whitaker model to calculate capillary pressure and liquid velocity, assumes that the liquid in the pores dries in order of decreasing pore size. If we add another assumption to this – namely, that the precipitate precipitates in the pores that are currently drying – then the volume fraction occupied by precipitate distribution by radius and pore size can be easily calculated. An example of such a distribution is shown in Figure 6.
@@
Salt deposition in pores causes changes in the pore volume distribution by size. These changes are particularly noticeable in the near-surface region of the sphere, where (in the examples discussed here) more salt is deposited. Note that these changes do not affect the results presented above, since the integration in formulas (18), (20) is carried out over capillaries filled with solution, and in them (according to the assumption made) there is no precipitated salt yet, therefore the distribution $dV_{p}/dr_{p}$ at these sizes has not yet changed. However, these changes should be taken into account when simulating re-impregnation of the same pellet with a different solution.
To assess the change in distribution $dV_{]}/dr_{p}$ during drying, one can subtract the distribution of precioitation volume by pore size from the initial distribution. An example of the result of such an operation is shown in Figure 7. This is only an approximate estimate, since it does not take into account the possible change in pore size when sediment falls into them. In addition, during subsequent steps to prepare the catalyst, the salt must turn into a catalytically active substance with a different volume.
Figure 7: Estimation of the pore volume distribution along the pore size in the near-surface layer of the sphere after drying. In large pores, no precipitate forms because the salt concentration in the solution is less than critical. The dashed line shows the original distribution.,
@@ 4 Conclusion
This paper is devoted to mathematical modeling of a complex multiparameter process of forming the spatial distribution of a catalyst precursor in a solution-impregnated porous sphere during the drying stage. The presented model is based on the Whitaker model of drying porous materials (ignoring the gas phase in the pores) and a model of precipitation from a supersaturated solution. The model allows for studying the influence of various factors on the final precipitate distribution. However, only qualitative agreement with experimental data can be expected, since the values of many process parameters are known only approximately. Also, little information is available on the dependence of the physical properties of the solution on temperature and salt concentration.
The drying model uses a representation of the porous medium as a bundle of capillaries of varying sizes. This representation establishes the order of pore drying; the solution in the pores dries in descending order of pore size. Assuming that the precipitate precipitates in the pores that are currently drying, the precipitate distribution by pore size can be estimated. The article provides an example of calculating the distribution of sediment by the radius of a sphere and the size of pores during drying of a sphere impregnated with a solution of ferrous sulfate.

### v2

@@ 3.4 Distribution of precipitation by pore size
The capillary bundle model, which is used in the Whitaker model to calculate capillary pressure and liquid velocity, assumes that the liquid in the pores dries in order of decreasing pore size. If one adds another assumption to this – namely, that the precipitation occurs in the pores that are currently drying – then the volume fraction occupied by precipitate distribution by radius and pore size can be easily calculated. An example of such a distribution is shown in Figure 6.
@@
Salt deposition in pores causes changes in the pore volume distribution by size. These changes are particularly noticeable in the near-surface region of the sphere, where (in the examples discussed here) more salt is deposited. Note that these changes do not affect the results presented above, since the integration in formulas (18), (20) is carried out over capillaries filled with liquid, and in them (according to the assumption made) there is no precipitated salt yet, therefore the distribution $dV_{p}/dr_{p}$ at these sizes has not yet changed. However, these changes should be taken into account when simulating re-impregnation of the same pellet with a different precursor solution.
To assess the change in distribution $dV_{]}/dr_{p}$ during drying, one can subtract the distribution of precipitation volume by pore size from the initial distribution. An example of the result of such an operation is shown in Figure 7. This is only an approximate estimate, since it does not take into account the possible change in pore size when salt falls into them. In addition, during subsequent steps to prepare the catalyst, the salt must turn into a catalytically active substance with a different volume.
Figure 7: Estimation of the pore volume distribution over the pore size in the near-surface layer of the sphere after drying. In large pores, no precipitate forms because the salt concentration in the solution is less than critical. The dashed line shows the original distribution.,
@@ 4 Conclusion
The above-presented mathematical model of drying an aqueous solution of a catalyst precursor in a porous sphere with precipitation of the dissolved substance in the pores is essentially a modified version of the model described in [7, 8]. The main modifications are as follows.
Firstly, in accordance with the goal of this work: to model the distribution over the volume of a sphere of precipitation falling from a liquid during drying, the presented model describes the drying process only in the presence of liquid in the pores, i.e. at $\epsilon_{l}(r,t)>0$. Due to the boundary condition (21), any solution of system (1)-(4) with the initial state $\epsilon_{l}(r,0)>0$ asymptotically tends to the state in which $\epsilon_{l}(r,t)=0$ while remaining positive, $\epsilon_{l}(r,t)\to+0$ at $t\to\infty$. Thus, in this model, liquid is present in the pores throughout the entire drying period. Dynamic equilibrium between liquid and water vapor is assumed, i.e., the water vapor density in the pores $\rho_{v}$ is always equal to the saturated vapor density $\rho_{v}^{*}$.
Secondly, since the value of $\rho_{v}^{*}$ under the conditions under consideration is more than three orders of magnitude less than the density of the liquid, the model does not take into account the transfer of mass and heat in the gas phase in the pores. It is assumed that the gas pressure in the pores is always equal to the pressure of the external gas $P_{gmex}$.
The Whitaker drying model uses a representation of the porous medium as a bundle of capillaries of varying sizes. This representation establishes the order of pore drying; the liquid in the pores dries in descending order of pore size. Assuming that the dissolved salt precipitates in the pores that are currently drying, the precipitate distribution by pore size can be estimated.
This paper demonstrates the influence of certain model parameters on the final distribution of precipitated salt across the volume of a sphere. The resulting distributions increase monotonically from the center to the surface of the sphere. A detailed parametric analysis of the model could reveal other distribution shapes It can be assumed that the shape of the final distribution is primarily determined by the hydrodynamics of the process, or more precisely, the liquid velocity distribution $v_{l}$. In this model, two factors influence liquid velocity: the surface tension of the liquid $\sigma_{l}$ and the pore volume distribution by size $dV_{p}/dr_{p}$, which were not varied in this study.
@@ Acknowledgments
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.